Please show me how to calculate the Slope and Intercept for this equation.
36X+9Y=18 also for this equation Y+3X+2. Thank you so very much.

Answer by checkley75(3666)

THE SLOPE (m) & THE Y INTERCEPT (b) IN THE EQUATION
Y=mX+b FOLLOWS:
36X+9Y=18
9Y=-36X+18
Y=-36/9+18/9
Y=-4X+2 THUS THE SLOPE (m)=-4 & THE Y INTERCEPT (b)=2
---------------------------------------------------------------------------
Y+3X+2=0
Y=-3X-2
HERE THE SLOPE (m)=-3 & THE Y INTERCEPT (b)=-2